Near the equator, the earth's magnetic field points almost horizontally to the north and has magnitude B = 0.5 x 10 ^-4 T. What should be the magnitude and direction for the velocity of an electron if its weight is exactly balanced by the magnetic pole? 


1
Expert's answer
2022-05-08T14:29:37-0400
"mg=qvB"

"v=\frac{mg}{qB}=\frac{9.1*10^{-31}*9.8}{1.6*10^{-19}*0.5*10^{-4}}=1.1*10^{-6}\:\rm m/s"

The velocity should be directed westward.